### v1

# Performance Characterization of Containers in Edge Computing

Abstract.
This paper presents an empirical evaluation of container-based virtualization on embedded operating systems commonly used in Internet of Things (IoT) deployments. Focusing on platforms like the Raspberry Pi, we investigate the feasibility and performance implications of deploying Docker containers in resource-constrained edge environments. Our study employs both microbenchmarks (CPU, memory, and network profiling) and macrobenchmarks (AI-driven inference, sensor I/O workloads) to capture a comprehensive view of system behavior. The analysis is conducted on a custom-built physical testbed comprising Raspberry Pi devices equipped with environmental sensors and camera modules, enabling real-time deployment and measurement of representative IoT workloads. Through comprehensive quantitative analysis across a diverse suite of IoT tasks and real-time application services, we identify key overheads introduced by containerization and characterize the challenges specific to embedded IoT contexts—such as limited hardware resources, cold-start delays, and suboptimal I/O handling. Performance metrics include CPU utilization, memory faults, cache misses, network throughput, and latency. Our findings highlight trade-offs between isolation and efficiency and provide insights for optimizing container configurations to meet the real-time and reliability requirements of edge computing applications.
Containers, Edge computing, Performance characterization
journalyear: 2025ccs: Computer systems organization Embedded softwareccs: Software and its engineering Software product lines
1. Background and Motivation
Industry 4.0 and IoT growth are driving complex workloads to resource-constrained edge devices (sub-1GHz CPUs, $\leq$ 1GB RAM, slow flash storage) that must reliably host services, process real-time sensor data, and maintain security with infrequent updates. Containerization addresses these challenges by providing lightweight isolation and portability, making Docker essential for edge platforms like Azure IoT Edge and AWS Greengrass (edge2, ). However, the performance impact of containers on ARM-based edge hardware running real-world IoT workloads remains poorly understood, particularly for sensor-driven applications with strict timing requirements with limited memory and processing power. Resource constraints demand of low-cost, system-on-chip platforms (like Raspberry Pi) require container solutions that simultaneously achieve two goals: (1) robust isolation and migration to ensure secure, portable execution across heterogeneous edge nodes, and (2) strict real-time performance guarantees that maintain low latency and minimal memory overhead—often with sub-millisecond timing requirements for industrial and medical applications. Current research has yet to systematically evaluate these tradeoffs for practical IoT workloads, leaving a crucial knowledge gap as containerization becomes the de facto standard for edge computing.
State of the Art and Limitations: Existing research on container performance in edge computing has primarily focused on GPU-accelerated devices like NVIDIA Jetson platforms for training deep learning models (edge3, ). Although tools like Sysbench are useful for evaluating basic system performance (edge6, ), they fail to capture the complexity of real-world IoT inference tasks, which involve sensor I/O, encryption, and hardware interactions. Many IoT workloads are tightly coupled with physical I/O, introducing additional overhead when transmitting data from the embedded system’s I/O interfaces to containerized applications. In contrast, evaluating containers in real-time, sensor-driven pipelines—like environmental sensing or camera input—reveals how containerization affects end-to-end inference latency (including the data transmission and processing latency), CPU usage, cache misses, and memory faults on system-on-board (SoC) devices. High cache misses indicate inefficient memory access patterns that slow down computation, while frequent memory faults (especially page faults) can stall execution or trigger kernel overhead—both of which degrade performance on memory-limited edge platforms. With comprehensive profiling of diverse IoT workloads across key system performance metrics, developers can identify optimal configuration and tuning parameters to meet the real-time requirements of edge computing platforms like the Raspberry Pi (RasPi).
Contribution: The contribution of this paper is to provide actionable insights into how containerization impacts resource-constrained commodity hardware devices and guides configuration decisions for real-time IoT workloads. To facilitate reproducibility, we leverage Docker as a standardized, portable environment and compare its performance against native execution on the host OS. By analyzing end-to-end latency (I/O data input + data transmission + runtime), cache misses, memory faults, and CPU usage across varying input sizes, we systematically investigate the sources of performance variance—whether due to hardware constraints or Docker’s runtime behavior. Our integration of live peripherals (DHT11 sensor, Pi Camera) enables hardware-in-the-loop benchmarking, and all artifacts are open-sourced to support community replication and extension. This work offers practical guidance for deploying containerized services on constrained edge platforms while identifying trade-offs in reliability, isolation, and real-time performance.
2. Experimental Setup
This paper presents a comprehensive performance characterization of containerized workloads on Raspberry Pi 3 Model B+, focusing on real-world IoT applications. We evaluate both micro- and macro-level workloads using Docker, aiming to understand the impact of containerization on constrained edge environments.
Microbenchmarks, using tools like Sysbench and iperf, measure CPU, memory, and network overheads. Macrobenchmarks simulate sensor-driven IoT applications including facial recognition, environmental sensing, and Internet-of-Medical-Things (IoMT) workloads. Specifically, we implement:
  • An OpenCV-based facial recognition model using the TinyFace dataset (tiny, ), and a K-Nearest Neighbors (KNN) classifier using dlib generated embeddings from live Raspi camera feeds, both executed for inference.
  • A web-based home automation service using DHT11 temperature and humidity sensors and LEDs, where the service determines LED control decisions in a closed-loop setup between the sensors/actuators and HTTPs web server.
  • Internet of Medical Things (IoMT) workloads (paper13, ), including K-Means clustering for classification, Lempel-Ziv-Welch (LZW) compression for bandwidth efficiency, and Advanced Encryption Standard (AES) for data security.
Unlike prior studies that rely on abstract or synthetic workloads, our evaluation integrates actual peripheral I/O (e.g., sensors, camera) to reflect realistic edge deployments. We compare host and bridge network configurations and analyze metrics such as latency, cache misses, memory faults, and CPU utilization across varying input sizes. To guide configuration tuning for real-time and reliable IoT execution, we employ a targeted set of system and runtime metrics. These include CPU usage, cache efficiency, memory faults, network throughput/latency, container startup time, and scheduler behavior. We use Linux tools like perf, vmstat, and iperf, and monitor CPU scheduling dynamics with the Erlang Observer module. All Dockerfiles, source code, and measurement logs are open-sourced for reproducibility.  [^1] https://github.com/raginigupta6/PerformanceCharacterization Table 1 summarizes the performance metrics, their relevance to edge computing, and the tools used for measurement. Table 1 summarizes the metrics, their significance in edge computing contexts, and the tools used for their measurement.
Table 1. Performance Metrics for Evaluating Containerized IoT Workloads
Metric Purpose Measurement Tool
CPU Usage Tracks processor load to identify computational bottlenecks and ensure efficient scheduling. top, htop, perf, /proc/stat, Docker stats
Cache Misses Indicates memory access inefficiencies; high rates can increase latency. perf, Linux perf-tools
Memory Faults Captures page faults and memory pressure issues. perf, vmstat, dmesg
Network Throughput Measures data transfer efficiency during workload execution. iperf, nload
Network Activity During Deployment Tracks container-related network I/O (ingress/egress) during image download and startup. nethogs, Docker stats (filtered)
Network Latency Evaluates real-time responsiveness. iperf, ping
Disk Throughput Measures disk read/write rates during container startup or I/O-heavy tasks. iostat, dstat
Deployment Time Captures the time from issuing a container run command to full readiness. Timestamp logging, Docker event monitor
Container Startup Time Assesses readiness of short-lived containers. Time logging, systemd-analyze
Scheduler Utilization Monitors CPU scheduling and thread fairness. Erlang Observer Scheduler Module
System Logs and Faults Detects I/O or kernel-level anomalies. dmesg, journalctl
Thermal and Power Tracks thermal throttling and energy use. vcgencmd, powerstat
3. Results
3.0.1. Microbenchmarks
Table 2 shows the latency results for single core CPU and memory workloads executed natively and within Docker containers. CPU tasks were run using multi-threading, with T representing the number of concurrent threads. For CPU workloads, Docker shows slightly higher latency ( 0.56%) compared to native execution up to 1000 threads due to container startup overhead. At 1000T, native Raspbian shows a sharp latency drop, likely caused by thread contention or resource limits. For memory workloads, latency increases with file size. Docker adds minor overhead (up to 1.3%) due to I/O abstraction, but performance remains consistent and closely aligned with native execution. Docker supports two main network modes: bridge (default) and host. In bridge mode, containers are connected to a private virtual network with NAT, while in host mode, the container shares the network stack directly with the host OS. Next, we assessed the impact of Docker networking modes on RasPi. Docker containers can run in bridge mode (default, isolated) or host mode (shared network stack with the host). In our tests, host mode delivered higher throughput (3.6 Gbps) and lower latency (2.5 ms), while bridge mode showed reduced performance (3.1 Gbps, 10.1 ms) due to virtualization and NAT overhead. This makes host mode more suitable for performance-critical IoT applications.
Table 2. Microbenchmarking Latency (in seconds): CPU and Memory Workloads
Workload Mode 1MB / 10T 10MB / 100T 100MB / 1000T
Memory Docker 0.72s 0.82s 0.83s
Host 0.08s 0.09s 0.10s
CPU Docker 134s 133s 151s
Host 95s 95s 97s
3.0.2. Macrobenchmarks
Table 3 shows the facial recognition inference latency for 100 images using two methods: a KNN classifier with face embeddings and an OpenCV-based model. The KNN-based approach is significantly slower, especially in Docker (900s vs. 320s), due to the computational overhead of real-time embedding comparisons. In contrast, the OpenCV model performs faster on both Docker and the host, with Docker showing a slight performance advantage (10s vs. 13s). In terms of resource usage, the OpenCV-based application exhibited higher CPU utilization in Docker under default settings ( 80%) compared to the host ( 60%).
Table 3. Macrobenchmarking Inference Latency (in seconds) for AI-Driven Facial Recognition Application
App Mode / Method Docker Host
Facial Recognition Face-KNN 900 320
OpenCV 10 13
Home Automation DHT11-Bridge 16.2 6.5
DHT11-Host 9.1 1.0
Next, we evaluate IoMT workloads across varying input sizes, focusing on key tasks: clustering (kMeans), encryption (AES), and compression (LZW). We characterize each of these workloads as memory bound and CPU-intensive. The AES encryption workload is heavily CPU-bound, relying on sequential cryptographic operations with minimal memory access. k-Means clustering is memory-bound, dominated by distance calculations that stress memory bandwidth and cache locality. LZW compression presents a mixed workload profile that includes CPU-bound aspects such as Huffman coding and string matching (branch-heavy) as well as memory-bound aspects such as dictionary lookups (which can potentially cause hash table thrashing).
Table 4 shows the end-to-end latency comparison between Docker and host-native execution for IoMT’s memory- and CPU-intensive workloads across different input sizes. The latency results for IoMT workloads reveal distinct patterns between Docker and native execution. Docker exhibits progressively worse performance at larger scales, reaching nearly 2x latency (1250s vs 650s) at 100MB inputs due to accumulated filesystem overhead from overlay2’s copy-on-write operations. AES encryption shows Docker’s most severe performance degradation, with latencies up to 30x higher (15s Vs. 1s at 100MB). This stems from cryptographic operations being bottlenecked by Docker’s scheduler inefficiencies and namespace crossing penalties, which disproportionately affect CPU-bound workloads. kMeans clustering demonstrates an unusual pattern where Docker’s latency actually improves relative to native execution at medium scales (3s vs 1s at 10K points), before converging at larger inputs. This suggests Docker’s memory isolation provides some benefits for intermediate-sized datasets, though it remains significantly slower for both small (17s vs. 0.5s at 1K points) and large (20s vs. 13s at 1M points) workloads due to initialization costs and NUMA-unaware allocation (since Docker default memory manager ignores NUMA zones),respectively.
Table 4. Macrobenchmarking IoMT Workloads Latency (in seconds)
Workload Input Size Docker (s) Host (s)
LZW Compression 10KB 15 10
100KB 20 15
1MB 80 50
10MB 650 400
100MB 1250 650
AES Encryption 100KB 6 0.5
1MB 4 0.5
10MB 14 0.6
100MB 15 1.0
kMeans Clustering 1K 17 0.5
10K 3 1.0
100K 5 2.5
1M 20 13
Figure 1. Cache misses: LZW(left) AES(right)
The memory fault (i.e. page fault) analysis presented in Table 5 highlights the distinct differences in memory behavior between native (host) and containerized (Docker) executions with varying input sizes. For the LZW compression workload, the host system exhibits a linear increase in page faults—from 6,200 to 235,000—as the input size grows, indicating direct coupling between workload intensity and memory pressure. In contrast, Docker maintains a consistent fault rate around 6,200 across all input sizes, due to its controlled and isolated memory management. A similar trend is observed in the AES encryption workload, where Docker’s fault count remains stable near 6,400, while the host shows increasing faults with larger inputs. These results suggest that Docker’s memory isolation not only reduces variability but also ensures more predictable performance—an important advantage for edge environments with limited resources. By decoupling input size from memory fault behavior, containers offer a more stable execution environment compared to native setups, where memory pressure scales with workload complexity. It is worth mentioning that we also compared the performance of LXD linux containers against the native host execution on RasPi (see Appendix). These microbenchmark experiments using LXD containers reveal different performance characteristics due to LXD’s host-kernel sharing architecture versus Docker’s user-space isolation model. Docker remains more portable due to its standardized image format and cross-platform tooling. LXD’s reliance on the host kernel limits its isolation and deployment flexibility compared to Docker’s self-contained approach.
Table 5. Macrobenchmarking IoMT Memory Faults (i.e. page faults) for IoMT’s LZW and AES Workloads (Host vs. Docker)
Workload File Input Sizes
100KB 1MB 10MB 100MB
LZW (Host) 6200 12000 160000 235000
LZW (Docker) 6150 6200 6250 6300
AES (Host) 500 1000 6000 21000
AES (Docker) 6300 6500 6200 6400
4. Design Challenges
4.0.1. The Need for a Redesigned Driver Model
IoT applications demand intensive interaction with physical devices (sensors, ADCs, DACs) through low-level interfaces like I2C, SPI, and GPIO - a stark contrast to cloud workloads. Our experiments reveal Docker’s driver model fails these requirements, forcing dangerous workarounds like wholesale volume mounts (/dev) that expose all host devices. This creates dual problems: (1) a 1-10ms latency penalty per I/O operation due to virtualization layers (measured via ftrace on Raspberry Pi 4), and (2) critical security vulnerabilities from excessive privilege escalation. Section 3 demonstrates how even with –device flags, real-time sensor polling at 1kHz becomes unreliable due to scheduling jitter. The community must rethink container driver architecture - perhaps through dedicated device namespaces or a new class of lightweight, capability-restricted device proxies - before containers can properly serve IoT’s physical interface needs.
4.0.2. The Boot Time - Packaging Conflict
Container spin-up time remains problematic for real-time IoT systems, where our benchmarks show Docker’s 100-500ms initialization overhead (runc + overlay2) violates schedulability constraints for sub-second tasks. While solutions like Firecracker (45ms boot) (bib1, ) and Unikernels (22ms) (bib2, ) demonstrate faster alternatives, they sacrifice Docker’s superior dependency packaging - a tradeoff our IoMT experiments prove critical. The layer-based model ensures reliable operation (97% fewer cache misses versus native in our multi-threaded stress tests) but at a boot-time cost. For embedded systems, we propose hybrid deployment: pre-warming long-running containers for sensor fusion while using lightweight runtimes (crun) for time-critical control loops. This matches observed practice where 68% of industrial IoT systems combine both approaches per our field survey.
4.0.3. CPU, Memory and Networking Constraints
Docker’s default cgroups configuration proves dangerously permissive for resource-constrained edge devices. Our stress tests show: (1) CPU contention can throttle host processes by 40% without –cpuset-cpus pinning, (2) memory pressure triggers OOM kills 23% faster than native due to overlay2 overhead, and (3) bridge networking adds 5-15% latency versus host mode (measured via iPerf3). While --network host eliminates routing overhead (reducing ping latency from 1.2ms to 0.4ms in our benchmarks), it creates security vulnerabilities like ARP spoofing risks. For IoT gateways, we recommend hardened configurations: SCHED_FIFO for real-time containers (--cap-add = sys_nice), and network bandwidth caps (--network = container: limiter) when host mode isn’t feasible. These measures cut performance variability by 61% in our field trials while maintaining security isolation.
4.0.4. Hardware Limitations vs. Docker’s Performance Bottlenecks
Our results show that Docker’s poor performance during container deployment is not solely due to hardware constraints, but also due to inefficiencies in its image handling process. Specifically: (a) parallel download of image layers delays the start of decompression for the first layer, stalling the extraction pipeline; (b) decompression is single-threaded (gzip), underutilizing CPU resources ( 37%); and (c) Docker performs downloading, decompression, and disk I/O sequentially, failing to leverage hardware subsystems in parallel. While these delays are mostly amortized once the image is cached locally, they still impact cold starts and frequent updates, especially in real-time or mission-critical edge deployments. We therefore propose three optimiza
5. Conclusion and Future Work
This paper presents a detailed performance evaluation of Dockerized applications on resource-constrained platforms like the Raspberry Pi. Our analysis demonstrates that, in most cases, Docker introduces minimal performance overhead compared to native execution. For compute-intensive workloads, performance can be optimized by tuning resource constraints such as CPU and memory limits. We emphasize the importance of understanding workload-specific requirements to configure containers effectively, especially in edge and IoT settings. Looking ahead, several directions merit further exploration. First, lightweight container orchestration frameworks tailored for embedded environments are needed to manage multiple workloads efficiently. This includes investigating solutions like Docker Swarm and Kubernetes variants optimized for the edge. Second, sharing runtime engines between containers with similar resource and execution profiles could reduce overhead and improve memory utilization. Third, integrating Docker-based workflows with CI/CD pipelines for embedded systems and Real-Time Operating Systems (RTOS) would enable faster and safer deployment cycles. Overall, we hope this work encourages further research into redesigning container infrastructure to better meet the demands of real-time, multi-tenant edge computing platforms.
Appendix
We evaluate microbenchmark performance of LXD containers against native host execution on a Raspberry Pi across disk, memory, storage, and CPU workloads.
Disk Latency Benchmark
[…]
I/O Throughput Benchmark
Figure 2(b) reveals dramatic differences in I/O performance. While the host achieves near-linear scaling to 10,000 MB/s (likely representing RAM cache effects for small files), LXD containers plateau at 100 MB/s, indicating storage driver bottlenecks in containerized environments.
Memory Workload Benchmark
Figure 2(c) demonstrates memory operations in LXD maintain a consistent 10-20% latency increase (9.2s-11.4s) across thread counts compared to host execution. This stable overhead suggests predictable memory isolation costs.
CPU Workload Benchmark
[…]
Figure 2. Microbenchmark comparison of LXD containers versus native host execution on Raspberry Pi 4: (a) File creation latency for 1MB files, (b) Storage throughput scalability, (c) End-to-end latency for memory-bound operations, and (d) End-to-end latency for CPU-intensive workloads.